# Circle bundles with PSC over some four manifolds

### v2

@@ 4.2. Proof of Corollary B
Let $X^{4}$ be a $T^{2}$-bundle over $T^{2}$. Then $M=X\#m\overline{\mathbf{CP}^{2}}$ satisfies the hypotheses required for Conclusions $(4)$ and $(5)$ of Theorem A.
Indeed, by a well-known observation of Thurston [Thu76], $X$ admits a symplectic structure. By the symplectic blowup construction due to Gompf [Gom95], $M$ also admits a symplectic structure. Note that $M$ is not aspherical for $m>1$ but admits a degree 1 map onto the aspherical manifold $T^{4}$. The fundamental group $\pi_{1}(X)$ is a $\mathbf{Z}^{2}$-extension over $\mathbf{Z}^{2}$, therefore it is amenable.
As a trivial case of the above, one could consider $X=T^{4}$. Note that here we also have $b_{2}^{+}=3>1$. Thus, $M$ satisfies the hypotheses required for Conclusions $(3),(4)$ and $(5)$.
@@
Let $(X^{2n},\omega)$ be any symplectically aspherical manifold with amenable fundamental group. By repeated applications of Theorem 1 and Proposition 2, we can produce a symplectic submanifold $M^{4}\subset X^{2n}$ such that $\pi_{1}(M)\cong\pi_{1}(X)$. Since $X$ is symplectically aspherical, so is $M$. Therefore, $M$ satisfies the hypotheses required for Conclusion $(5)$ in Theorem A.
As an explicit example, one could take $X=T^{6}$. Considering the symplectic structure on $T^{6}$ as the product $T^{4}\times T^{2}$, we see that the projection $M\to T^{4}$ to the first factor must be a positive degree map by the same argument as in Proposition 5. Therefore, $M$ satisfies the hypotheses of Conclusions $(3),(4)$ and $(5)$. One salient feature of this example is that $\pi_{1}(M)\cong\mathbf{Z}^{6}$ has cohomological dimension $6$. Therefore, the techniques in [BD16] are unable to provide apriori bounds on the macroscopic dimension of $E_{Z}$.
@@
(Examples within examples) Take any manifold $M$ satisfying the hypotheses required for Conclusions $(3),(4)$ and $(5)$ in Theorem A. Let $Z\subset M\times S^{2}$ be a Donaldson divisor as in the statement of Theorem A. Then $Z$ also satisfies the same hypotheses.
Proof of Corollary B.
Let $M$ be any manifold from Example 8, 9 or 10 above. Let $k\gg 1$ be sufficiently large. Let $Z_{k}\subset M\times S^{2}$ be the submanifold furnished by Donaldson’s theorem (Theorem 1), such that $[Z_{k}]=\mathrm{PD}(k\cdot[\omega])$. Let $L$ be the pre-quantum line bundle over $M$, so that $c_{1}(L)=[\omega]$. Let $E_{k}$ be the unit sphere bundle of $L^{\otimes k}$ restricted to $Z_{k}$. It follows from the proof of Theorem A that $Z_{k}$ and $E_{k}$ satisfy Conclusions $(1)$ and $(2)$. Proposition 4 implies $|\chi(Z_{k})|\to\infty$ as $k\to\infty$, as desired. ∎
4.3. Proof of Corollary C
We begin by recalling the definition of enlargeable manifolds from Gromov-Lawson [GL83].
Definition 4.
The basic example of an enlargeable manifold is the unit circle $S^{1}\cong\mathbf{R}/2\pi\mathbf{Z}$. Indeed, the $n$-sheeted cover of $\mathbf{R}/2\pi\mathbf{Z}$ is $\mathbf{R}/2\pi n\mathbf{Z}$. The map $\mathbf{R}/2\pi\mathbf{Z}\to\mathbf{R}/2\pi n\mathbf{Z}$ given as scaling by $1/n$ is a $1/n$-contracting map, for any $n\geq 1$. Also, $S^{1}$ is certainly a spin manifold. It is straightforward to see that
  1. (1)
    Product of enlargeable manifolds is enlargeable,
  2. (2)
Proof of Corollary C.
Let $M=T^{4}$ as in Example 8 with $m=0$. Let $\omega$ denote the product symplectic form on $T^{4}\times S^{2}$. Let $k\gg 1$ be a sufficiently large even integer. Then, by Theorem A, there exists a submanifold $Z^{4}\subset(T^{4}\times S^{2},\omega)$ such that $[Z]=\mathrm{PD}(k\cdot[\omega])$, $Z$ admits a positive degree map to $T^{4}$, and there exists a nontrivial circle bundle $E_{Z}\to Z$ such that the total space $E_{Z}$ admits a metric of positive scalar curvature.
Since $T^{4}$ is enlargeable and $Z$ already admits a positive degree map to $T^{4}$, to show that $Z$ is also enlargeable we just need to show that it is also spin, i.e. the second Stiefel-Whitney class is 00. Note that as $Z$ is a symplectic manifold, the second Stiefel-Whitney class of $Z$ is the first Chern class of $Z$ modulo $2$:
$w_{2}(Z)\equiv c_{1}(Z)\pmod{2}$
By Equation (2.3) in Proposition 4, we have
$c_{1}(Z)=j^{*}c_{1}(T^{4}\times S^{2})-k\cdot j^{*}[\omega]$
Next, we have $c_{1}(T^{4}\times S^{2})\equiv w_{1}(T^{4}\times S^{2})\equiv 0\pmod{2}$, as $T^{4}$ and $S^{2}$ are both spin, and product of spin manifolds is also spin. Since $k$ was moreover chosen to be even, we conclude $c_{1}(Z)\equiv 0\pmod{2}$. Thus, $w_{2}(Z)=0$. Hence, $Z$ is spin. ∎

### v3

@@ 4. Further examples and remarks
Let $m\geq 2$, and $Z^{2m}\subset(T^{2n},\omega)$ be a symplectic submanifold obtained by repeated applications of Theorem 1. Hence, by Proposition 2, $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{mc}(\widetilde{E})\leq n+1$. From the homotopy long exact sequence for fiber bundles, we obtain an exact sequence
$0\cong\pi_{2}(E)\to\pi_{1}(S^{1})\cong\mathbf{Z}\to\pi_{1}(E)\to\pi_{1}(Z)\to 0$
Therefore, $\pi_{1}(E)$ is an infinite cyclic extension of $\pi_{1}(Z)\cong\mathbf{Z}^{2n}$. Therefore, the asymptotic dimension of $\pi_{1}(E)$ is $\mathrm{asdim}(\pi_{1}(E))=2n+1$. Likewise, for the virtual cohomological dimension, we have $\mathrm{vcd}(\pi_{1}(E))=2n+1$. Note that for $m=n$, we get $\mathrm{vcd}(\pi_{1}(E))=\dim E$. Thus, in this case the techniques of [BD16] are unable to provide apriori upper bounds on $\dim_{mc}(E)$, even if $E$ is totally non-spin.
Next, we give examples of symplectic manifolds which are enlargeable. These can be used as an input in Part $(4)$ of Theorem A. To construct these examples, we use the recent work of Fine and Panov [FP21], which shows that every even dimensional closed oriented manifold is dominated by a symplectic manifold.
Example 11 (Enlargeable symplectic manifolds).
Let $N^{2n}$ be an even dimensional enlargeable manifold, for $n\geq 3$. By [FP21, Theorem 1], there exists a symplectic manifold $(M^{2n},\omega)$ and a positive degree map $f:M\to N$. In fact, we may arrange $c_{1}(M)$ to be even by ensuring that in the proof of [FP21, Theorem 1], the chain of symplectic submanifolds of the twistor space is selected so that each of them is a Donaldson divisor that is Poincaré dual to an even (respectively, odd) multiple of the symplectic form in the previous submanifold, depending on if the parity of $n$ is even (respectively, odd). This ensures $c_{1}(M)$ is even by the remark at the end of the article [FP21, pg. 3] regarding the first Chern class of the twistor space. Therefore, we have
$w_{2}(M)\equiv c_{1}(M)\equiv 0\pmod{2}$
Thus, $M$ is spin. Since $M$ admits a positive degree map to the enlargeable manifold $N$, $M$ must itself be enlargeable. Therefore, for any surface with an area form $(S_{g},\omega_{0})$, the symplectic manifold $(M\times S_{g},\omega\oplus\omega_{0})$ satisfies the hypothesis of Part $(4)$ of Theorem A.
We now give some examples which do not satisfy the hypotheses of Parts (3) and (4) in Theorem A, nonetheless we observe that they satisfy the conclusions using similar methods.
Example 12 (Product with a $2$-sphere).
Let $(M^{2n},\omega)$ be an integral symplectic manifold which is symplectically aspherical such that $\pi_{1}(M)$ is amenable. Let $\omega_{0}$ denote the standard area form on $S^{2}$. Then $(M\times S^{2},\omega\oplus\omega_{0})$ is not symplectically aspherical, hence it fails the hypothesis of Part $(3)$ of Theorem A.
Nevertheless, let $Z^{2n}\subset(M\times S^{2},\omega\oplus\omega_{0})$ be a Donaldson divisor. Consider the map $f:M\times S^{2}\to M$ given by the projection to the first factor. By Proposition 5, $f|_{Z}:Z\to M$ is a positive degree map. Since $M$ is symplectically aspherical, it is rationally essential due to Proposition 7. Since $f|_{Z}$ has positive degree, $Z$ must also be rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)$. Therefore, $Z$ is a rationally essential manifold with amenable fundamental group. By [Dra11, Theorem 7.2], $\dim_{mc}(\widetilde{Z})=2n$.
Example 13 (Symplectic blow-up).
Let $M=T^{2n}\#\overline{\mathbf{CP}}^{n}$ for $n\geq 3$. By the symplectic blowup construction due to Gompf [Gom95], $M$ admits a symplectic structure $\omega$. Let $\overline{\mathbf{CP}}^{1}$ denote the exceptional divisor. Then, $\overline{\mathbf{CP}}^{1}\subset(M,\omega)$ is a symplectic submanifold. Therefore,
$\int_{\overline{\mathbf{CP}}^{1}}\omega\neq 0$
Thus, $(M,\omega)$ is not symplectically aspherical. In addition, $M$ is not a product of an enlargeable manifold with a surface. Therefore, $(M,\omega)$ does not satisfy the hypothesis of either Part $(3)$ or Part $(4)$ in Theorem A.
Nevertheless, let $Z^{2n-2}\subset(M,\omega)$ be a Donaldson divisor with $[Z]=\mathrm{PD}(k[\omega])$, for some $k\gg 1$. Consider the map $f:M\to T^{2n-2}$ given by composing the blow-down map $q:M\to M/\overline{\mathbf{CP}}^{1}\cong T^{2n}$ with the projection $T^{2n}\cong T^{2n-2}\times T^{2}\to T^{2n-2}$. For $z\in T^{2n-2}$,
$(f|_{Z})^{-1}(z)=f^{-1}(z)\cap Z=q^{-1}(\{z\}\times T^{2})\cap Z$
Choose $z\in T^{2n-2}$ such that $\{z\}\times T^{2}\subset T^{2n}$ is disjoint from a neighborhood of the blown up point $\{z_{0}\}=q(\overline{\mathbf{CP}}^{n})\subset T^{2n}$. By the construction of symplectic blow-ups [Gom95], $q$ restricts to a symplectomorphism on the complement of a neighborhood of the exceptional divisor $\overline{\mathbf{CP}}^{1}\subset M$. Therefore, $T^{2}\cong q^{-1}(\{z\}\times T^{2})\subset(M,\omega)$ is a symplectic submanifold intersecting $Z^{2n-2}$ transversely. Therefore, the oriented intersection number is given by:
#$\#(q^{-1}(\{z\}\times T^{2})\cap Z)=k\int_{q^{-1}(\{z\}\times T^{2})}\omega>0$
Thus, $f|_{Z}:Z\to T^{2n-2}$ is a map of positive degree. Hence, $Z$ is rationally essential. By Proposition 2, $\pi_{1}(Z)\cong\pi_{1}(M)\cong\mathbf{Z}^{2n}$ which is amenable. Therefore, by [Dra11, Theorem 7.2], $\dim_{mc}(\widetilde{Z})=2n-2$. Additionally, if $n=3$ (resp. $n=4$), then $Z$ admits a positive degree map to $T^{4}$ (resp. $T^{6}$). In that case, $Z$ is not PSC. [SY79]
In dimension $4$, it is possible to produce many more examples of non-PSC manifolds admitting circle bundles with total space PSC using a deep result of Taubes [Tau94].
@@
Let $(M^{4},\omega)$ be an integral symplectic $4$-manifold with $b_{2}^{+}(M)>1$, and let $\omega_{0}$ denote an integral area form on the surface $S_{g}$. Let $Z^{4}\subset(M\times S_{g},\omega\oplus\omega_{0})$ be a Donaldson divisor. By Proposition 5, the projection $f:Z\to M$ has positive degree. Therefore, the induced map
$f^{*}:H^{*}(M;\mathbf{Q})\to H^{*}(Z;\mathbf{Q})$
is an injection of quadratic spaces, where both the domain and range are equipped with the quadratic form given by cup product. Therefore, since $b_{2}^{+}(M)>1$, we must have $b_{2}^{+}(Z)>1$ as well. Since $Z$ is a symplectic $4$-manifold with $b_{2}^{+}(Z)>1$, by Taubes’ theorem [Tau94], it follows that $Z$ has non-vanishing Seiberg-Witten invariants. Therefore, $Z$ is not PSC. Nevertheless, by Part $(1)$ of Theorem A, $Z$ admits a circle bundle with total space being PSC.
We close this section with an example which illustrates that macroscopic dimensions of manifolds having diffeomorphic universal covers need not be the same. For more surprising examples, we refer the reader to Kastenholz-Reinhold [KR23].
@@
Let $M=T^{2n}$ for $n\geq 3$ and $\omega$ be the standard symplectic form on the torus, normalized to be integral. Let $Z\subset(M,\omega)$ be a Donaldson divisor with $[Z]=\mathrm{PD}(k[\omega])$. By Part $(2)$ of Theorem A, there exists a circle bundle $S^{1}\hookrightarrow E\stackrel{{\scriptstyle p}}{{\to}}Z$ such that $\dim_{mc}(\widetilde{E})\leq n$. Let $\widehat{E}\to\widetilde{Z}$ denote the circle bundle obtained by pulling back the bundle $p:E\to Z$ along the universal cover $q:\widetilde{Z}\to Z$. The resulting bundle is classified by an element in
$H^{2}(\widetilde{Z};\mathbf{Z})\cong\mathrm{Hom}(H_{2}(\widetilde{Z};\mathbf{Z}),\mathbf{Z})\cong\mathrm{Hom}(\pi_{2}(\widetilde{Z}),\mathbf{Z})\cong\mathrm{Hom}(\pi_{2}(Z),\mathbf{Z})$
More precisely, it is classified by the cohomology class of the $2$-form $q^{*}\omega$. However, since $(Z,\omega)$ is symplectically aspherical by Example 6, for any map $f:S^{2}\to\widetilde{Z}$,
$\int_{S^{2}}f^{*}(q^{*}\omega)=\int_{S^{2}}(q\circ f)^{*}\omega=0$
Therefore, $[q^{*}\omega]=0\in H^{2}(\widetilde{Z};\mathbf{Z})\cong\mathrm{Hom}(\pi_{2}(Z),\mathbf{Z})$. This implies that the bundle $\widehat{E}\to\widetilde{Z}$ is trivial, i.e., $\widehat{E}\cong\widetilde{Z}\times S^{1}$. Since the total space of $\widehat{E}$ is a cover of the total space of $E$, we obtain that $\widetilde{Z}\times S^{1}$ covers the total space of $E$. In particular, $\widetilde{E}$ is diffeomorphic to $\widetilde{Z}\times\mathbf{R}$. Note that $\dim_{mc}(\widetilde{Z})=2n-2$, by Part $(3)$ of Theorem A. Thus, $E$ and $Z\times S^{1}$ have diffeomorphic universal covers, but nevertheless have different macroscopic dimensions.